40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1686     Entry time: Tue Jul 26 11:57:45 2016
Author: awade and Antonio 
Type: Notes 
Category: North Cavity 
Subject: North cavity temperature changed at 11:55 

Tue Jul 26 11:55:00 2016, north refcavity heater voltage was changed from 10.4 V to 10.6 V. We will leave it to stabilise while we are at group meeting and lunch.

ELOG V3.1.3-